OpenStudy (arihangdu):

Iron - 59 has a half life of 45.1 ayes. How old is an iron nail if the Fe -59 content is 25% that of a new sample of iron ? Please show me all the calculation to understand.

OpenStudy (arihangdu):

Please help me I will give medals and fans.

Vocaloid (vocaloid):

amount left = original amount*(1/2)^(total time/length of half-life) so plugging in our numbers, we get: .25 = (1/2)^(t/45.1) solve for t

OpenStudy (arihangdu):

How to solve please help me

Vocaloid (vocaloid):

.25 = (1/2)^2 so (1/2)^2 = (1/2)^(t/45.1) 2 = t/45.1 solve for t

OpenStudy (arihangdu):

T= 45.1 x 2= 90?

OpenStudy (arihangdu):

Right?

Vocaloid (vocaloid):

90.2 but yeah

OpenStudy (arihangdu):

Any formula we need to use here please?

Vocaloid (vocaloid):

no, that's it, t = 90.2 years...

OpenStudy (arihangdu):

Days or years ?

OpenStudy (arihangdu):

A=Ao *m this not a decays formula ?or I confused !,

OpenStudy (arihangdu):

Ok thank you so much for helping

OpenStudy (arihangdu):

Please receive my appreciation.

OpenStudy (arihangdu):

Thanks for helping me.

OpenStudy (arihangdu):

I don't know how you do this steps .25= 1/2^t/45.1 1/2^2=1/2)^t/45.1